1. Suppose 21 million people are unemployed today in an economy. If the rate of job separation is 0.02 per month and the rate of job finding is 0.08 per month, what is the natural rate of unemployment? If labor force remains unchanged over time and 64 million people are employed in the steady-state of the economy, how many people are unemployed in the steady-state?
Answer #1

Natural unemployment rate = Job separation rate / (Job separation rate + Job finding rate)

= 0.02 / (0.02 + 0.08)

= 0.02 / 0.1

= 0.02 = 20%

Number employed = Labor force (LF) x Employment rate = LF x (1 - Unemployment rate) = LF x (1 - 0.2) = LF x 0.8

LF x 0.8 = 64 million

LF = 80 million

Number unemployed = LF x Unemployment rate = 80 million x 20% = 16 million